# Analog Spiking Neuron in CMOS 28 nm Towards Large-Scale Neuromorphic Processors

### v1

@@ Abstract
The computational complexity of deep learning algorithms has given rise to significant speed and memory challenges for the execution hardware. In energy-limited portable devices, highly efficient processing platforms are indispensable for reproducing the prowess afforded by much bulkier processing platforms.
@@
Index Terms:
Neuromorphic circuit, spiking analog CMOS neuron, machine learning, spiking neural networks (SNN), TinyML, EdgeAI.
@@ I Introduction
Biology-inspired neuron cells are the fundamental computing units of neuromorphic systems [1], and their energy efficiency is closely linked to them, highlighting the importance of the neuron model choice. In 1943, McCulloch and Pitts [2] proposed the first artificial neuron model, based on binary logic principles. They were followed in 1952 by Hodgkin and Huxley who developed a continuous, conductance-based, model that more accurately described the biological neuron activity [3].
@@
The McCulloch and Pitts model was closely linked to the digital computing paradigm with 1’s and 0’s, rendering it easier to simulate on digital computers than the differential equation-based model proposed by Hodgkin and Huxley, but the inter-neural connection weights of the model were manually set, severely limiting its learning capacity. In the sixties, Mead noticed that the differential equations governing the dynamics of biological neurons bore a close resemblance to those governing the subthreshold dynamics of CMOS transistors [4]. Therefore, a few transistors operating in the analog domain could replace thousands of transistors working collectively on a digital chip to emulate differential equations.
@@
Artificial spiking neurons aim to emulate the characteristics of biological neurons while maximizing the computational capabilities, flexibility and robustness of the neuromorphic systems that use them. Spiking neurons differ from the more common rate-coded neurons by their spiking instead of continuous output, which leads to substantial energy saving. The mathematical models are typically implemented using CMOS transistors, with each functionality incurring costs in terms of silicon footprint and power dissipation [5]. The key performance metrics of the resulting CMOS circuits consist of several parameters such as total energy consumption, chip area and biological plausibility.
@@
Energy consumption is estimated both in terms of static power consumption and dynamic energy per spike. Additional metrics include supply voltage, mean and maximum spiking frequency, communication support (AER, SPI, I2C), spiking and bursting patterns, runtime reconfigurability (thresholds, time constants, refractory period), spike frequency adaptation, intrinsic plasticity, load driving capability, synaptic integration with non-volatile memory (NVM) devices such as memristors, and robustness against Process Variation and Temperature (PVT) variations. Achieving an appropriate balance between these parameters for the target application is crucial for constructing efficient neuromorphic systems.
@@
The system of differential equations in the Hodgkin and Huxley model of a biophysical neuron is difficult to simulate on a large scale. In 2003, Izhikevich proposed a more efficient model in this respect, demonstrating an array of cortical spiking and bursting behaviors [6, 7]. Early silicon implementations of the model include the works in [8, 9, 10]. More recent work includes an implementation predicated on a ring oscillator in 65 nm CMOS node [11]. The spiking and bursting modes of this neuron [11] were dictated by digital control signals, which can be supplied by an FPGA. The circuit was optimized with respect to area, energy consumption, and display resilience against PVT variations as corroborated in [12].
@@
In yet another work [14], two neuron implementations are explained. A bio-inspired version, equipped with leak conductance, feedback loop, and related features, is proposed alongside a streamlined version that employs fewer transistors and enables a higher spiking rate. The energy expenditure per spike was evaluated at 4fJ/spike, with the spatial footprint of the simplified version determined to be 35 $\mu$m2(65nm). Subsequently, the proposed bio-inspired version was optimized on 55nm CMOS node in [15]. The first implementation, consuming a mere 2.6 fJ/spike, exhibited Fast Spiking, while the Low Threshold Spiking (LTS) second neuron with additional components, demonstrated spike frequency adaptation.
@@
The adaptive exponential Integrate and Fire neuron, also known as I&F neuron, includes spike frequency adaptation. Essentially, the neuron adjusts to repeated stimuli, leading to a rise in the voltage threshold and a subsequent decrease in the neuron’s firing rate. Boahen demonstrated the feasibility of spike-frequency adaptation by linking a four-transistor ”current-mirror integrator” in negative-feedback mode to any Integrate and Fire circuit [17]. This idea was further adopted by Indeveri et al. in [18], where the implementation of AdexIF neuron was showcased. This neuron concept underwent evolution over time, with feature size reduction and the incorporation of attributes such as threshold voltage and refractory period modulation. One implementation, featured in [19], employs 22 nm FDSOI technology. The AdeXIF neuron is incorporated in the DYNAPSE chip [20], representing one of the cutting-edge neuromorphic chips.
The leaky-integrate-and-fire modeling approach offers a fair compromise between computational complexity and biological features. The LIF is the simplest model to emulate a biological neuron, as it only accounts for the leaky integration and the spiking behaviour. Although many works have been published about modeling the spiking behavior of biological neurons, the LIF model is a generally accepted trade-off between the complexity of biologically plausible neuron models such as Hodgkin-Huxley and Izhikevich models [21, 22]. The simplicity of this model makes it promising for the implementation of large-scale NeuroSoC processors in real-life applications, as is the case for the IBM TrueNorth chip [23] and Intel Loihi [24]. However, they are implemented with digital architectures.
@@ II-A LIF Model
This equation represents a balance between two forces: the leaky term $-(V_{\text{mem}}(t)-V_{\text{reset}})$, which drives the membrane potential towards the reset potential, and the input current term $R_{m}I_{\text{syn}}(t)$, which pushes the membrane potential in the opposite direction.
@@
  • $dt$: the discrete time interval to update the membrane potential of the model. Smaller values result in more accurate simulations at the cost of increased computational complexity.
@@ II-B LIF Model Computation Sequence
for each time step $dt$ do
@@ II-C LIF Implementation in CMOS
The Spike generation block comprises two back-to-back CMOS inverters that generate artificial action potentials at the neuron’s output. The inverters consist of transistors $M_{5}$, $M_{6}$, $M_{7}$, and $M_{8}$. When $V_{mem}$ reaches the inversion threshold $V_{\text{inv}}$, the inverters are activated producing a spike at the output. After producing a spike, the Reset block is activated to discharge the membrane capacitor $C_{mem}$ and prepare the neuron for the next cycle. Where the inversion voltage $V_{\text{inv}}$ is expressed as [25] :
$V_{inv}=\frac{V_{DD}}{2}+\frac{1}{2}\left(V_{thn}-|V_{thp}|\right)+n_{n}V_{T}\ln\left[\sqrt{\frac{I_{0p}}{I_{0n}}}\right]$ (5)
@@ II-C1 Sub-Threshold Operation
Sub-threshold operation in MOS transistors occurs when the gate-to-source voltage ($V_{GS}$) is below the threshold voltage ($V_{th}$), resulting in an exponential but weak increase in the drain current ($I_{D}$) with the slight change in $V_{GS}$ [26]. Hence, the drain current requires careful analysis to understand its impact on circuit behavior. In this particular mode, the drain current of a MOS transistor is given by [25]:
$I_{D}=(n-1)I_{0}e^{\frac{q(V_{GS}-V_{th})}{nk_{B}T}}\left(1-e^{-\frac{qV_{DS}}{k_{B}T}}\right)$ (6)
Where:
$n=1+\frac{C_{\Sigma}}{C_{ox}}$ (7)
$C_{\Sigma}$ being the depletion region capacitance and $I_{0}$ given by:
$I_{0}=\mu_{0}C_{ox}\frac{W}{L}\left(n-1\right)V_{th}^{2}$ (8)
As we typically have $\left(1-e^{-\frac{qV_{DS}}{k_{B}T}}\right)\approx 1$, equation 6 can be approximated by:
$I_{D}=\mu_{0}C_{ox}\frac{W}{L}\left(\frac{C_{\Sigma}}{C_{ox}}\right)^{2}V_{th}^{2}\;e^{\left(\frac{q(V_{GS}-V_{th})}{nk_{B}T}\right)}$ (9)
Eq. (9) showcases the exponential dependency of the drain current to the gate-source voltage which is highly counter-intuitive.
Let’s apply the latter equation to the input currents $I_{syn}$ and $I_{mem}$ in Fig. 1. The source current can be modeled as a resistor since our input characterization signals are DC. Thus, $V_{SG1}=V_{DD}-R_{syn}I_{syn}$ where $R_{syn}$ is the equivalent impedance seen from the drain of $M_{1}$.
and $I_{mem}$ is given by:
$I_{\text{mem}}=I_{0P}\;e^{\left(\frac{V_{SG2}-V_{th}}{nV_{T}}\right)}$ (10)
On the other hand $V_{SG1}=V_{SG2}$:
$I_{\text{mem}}=I_{0P}\;e^{\left(\frac{V_{DD}-R_{\text{syn}}I_{\text{syn}}-V_{th}}{nV_{T}}\right)}$ (11)
Replacing in Eq. 4 while neglecting $V_{reset}$ yields:
$V^{\text{max}}_{\text{mem}}(T)=\frac{1}{C_{\text{mem}}}\cdot I_{0P}T\cdot e^{\left(\frac{V_{DD}-R_{\text{syn}}I_{\text{syn}}-V_{th}}{nV_{T}}\right)}$ (12)
Which States that $Vmem$ peak’s voltage is dependent on the input stimulation current $I_{syn}$.
@@
Given the spike generation bloc of cascaded CMOS inverters $M_{5}$-$M_{6}$ and $M_{7}$-$M_{8}$, it is worth noting that the output-high voltage or $V_{OH}$ across $M_{8}$ is not necessarily equal to VDD due to the sub-threshold mode of operation [25]. We have:
$V_{OH}=V_{DD}\\ -V_{T}\ln{\left(\frac{1}{1-\frac{I_{0N}}{I_{0P}}e^{\left(\frac{-V_{DD}-V_{thn}+\mid V_{thp}\mid}{n_{n}V_{T}}\right)}}\right)}$ (13)
Where $V_{T}=\frac{k_{B}T}{q}$ is the thermal voltage approximately equal to $26mV$ at room temperature. The natural logarithmic part of the equation is negative therefore Eq. (13) yields that $V_{OH}$ is lower than $V_{DD}$. Consequently, the previous equation predicts some variability in the spike’s output voltage amplitude and the variance of the up-swing triggering in the membrane voltage.
Propagation Delay
The propagation delay of inverter $M_{5}$-$M_{6}$, $t_{P}$, is defined as the average of the low-to-high and the high-to-low propagation delays, $t_{PLH}$ and $t_{PHL}$, respectively [27]. $t_{PHL}$ can be found by averaging the discharging current of $C_{out}$ from [25]:
$t_{PHL}=\frac{(C_{gs7}+C_{gs8}+C_{gd7}+C_{gd8})V_{OH}}{I_{D6}}$ (14)
Where $C_{gs}$ is the gate-source capacitance and $C_{gd}$ the gate-drain capacitance given, respectively, by the following expressions [28]:
$C_{gs}=\frac{W}{L}C_{ox}\frac{2(1+2\eta)}{3(1+\eta)^{2}}$ (15)
$C_{total}=\frac{W}{L}C_{ox}\frac{2(\eta^{2}+2\eta)}{3(1+\eta)^{2}}$ (16)
Where $\eta$ is the drain-induced barrier lowering (DIBL) coefficient which represents how deep a transistor is in sub-threshold or how much it is not in saturation. Typical values for $\eta$ are 0.0266 and 0.0351 for the NMOS and PMOS devices [25].
Given the propagation delay dependence on the drain current and the exponential dependence of the drain current on the gate-source voltage in inverter $M_{5}$-$M_{6}$, the gate-source voltage, which is in this case the membrane voltage $V_{mem}$, will have a dissimilar effect on the inverter since the slope of $V_{mem}$ is equal to $I_{mem}/C_{mem}$.
This leads to the membrane voltage triggering at different levels because $I_{D5}$ varies exponentially hence activating the $M_{3}$ reset transistor, resulting in an up-swing of the membrane voltage up to VDD [5], at different $V_{mem}$ levels depending on it’s slope steepness.

### v2

# A 1.6-fJ/Spike Subthreshold Analog Spiking Neuron in 28 nm CMOS

@@ Abstract
The computational complexity of deep learning algorithms has given rise to significant speed and memory challenges for the execution hardware. In energy-limited portable devices, highly efficient processing platforms are indispensable for reproducing the efficiency achieved by much bulkier platforms.
@@ I Introduction
Biology-inspired neuron cells are the fundamental computing units of neuromorphic systems [3], and their energy efficiency is closely linked to them, highlighting the importance of the neuron model choice. In 1943, [22] proposed the first artificial neuron model, based on binary logic principles. They were followed in 1952 by [12] who developed a continuous, conductance-based, model that more accurately described the biological neuron activity.
@@
[22] model was closely linked to the digital computing paradigm with 1’s and 0’s, rendering it easier to simulate on digital computers than the differential equation-based model proposed by [12], but the inter-neural connection weights of the model were manually set, severely limiting its learning capacity. In the sixties, Mead noticed that the differential equations governing the dynamics of biological neurons bore a close resemblance to those governing the dynamics of CMOS transistors [19]. Therefore, a few transistors operating in the analog domain could replace thousands of transistors working collectively on a digital chip to emulate differential equations.
@@
Artificial spiking neurons aim to emulate the characteristics of biological neurons while maximizing the computational capabilities, flexibility and robustness of the neuromorphic systems that use them. Spiking neurons differ from the former by their intermittent spiking feature instead of continuous output, which leads to substantial energy saving. The mathematical models are typically implemented using CMOS transistors, with each functionality incurring costs in terms of silicon footprint and power dissipation [15]. The key performance metrics of the resulting CMOS circuits consist of several parameters such as total energy consumption, chip area and biological plausibility.
@@
Energy consumption is estimated both in terms of static power consumption and dynamic energy per spike. Additional metrics include supply voltage, mean and maximum spiking frequency, spiking and bursting patterns, runtime reconfigurability (thresholds, time constants, refractory period), spike frequency adaptation, intrinsic plasticity, load driving capability, synaptic integration with non-volatile memory (NVM) devices such as memristors, and robustness against Process Variation and Temperature (PVT) variations. Achieving the right balance of these parameters is crucial for building efficient neuromorphic systems for the target application.
@@
The system of differential equations in the Hodgkin and Huxley model of a biophysical neuron is difficult to simulate on a large scale. In 2003, [16, 17] proposed a more efficient model in this respect, demonstrating an array of cortical spiking and bursting behaviors [16, 17]. Early silicon implementations of the model include the works in [33, 24, 30]. More recent work includes an implementation predicated on a ring oscillator in 65 nm CMOS node [37]. The spiking and bursting modes of this neuron [37] were dictated by digital control signals, which can be supplied by an FPGA. The circuit was optimized with respect to area, energy consumption, and display resilience against PVT variations as corroborated in [38].
@@
[…]
@@
The adaptive exponential Integrate and Fire neuron, also known as I&F neuron, includes spike frequency adaptation. Essentially, the neuron adjusts to repeated stimuli, leading to a rise in the voltage threshold and a subsequent decrease in the neuron’s firing rate. [4] demonstrated the feasibility of spike-frequency adaptation by linking a four-transistor ”current-mirror integrator” in negative-feedback mode to any Integrate and Fire circuit. This idea was further adopted by [14], where the implementation of AdexIF neuron was showcased.
This neuron concept underwent evolution over time, with feature size reduction and the incorporation of attributes such as threshold voltage and refractory period modulation. One implementation is featured in [26] and is fabricated in 22 nm CMOS FDSOI technology.
The leaky-integrate-and-fire modeling approach offers a fair compromise between computational complexity and biological features. The LIF is the simplest model to emulate a biological neuron, as it only accounts for the leaky integration and the spiking behaviour. Although many works have been published about modeling the spiking behavior of biological neurons, the LIF model is a generally accepted trade-off between the complexity of biologically plausible neuron models [11]. The simplicity of this model makes it promising for the implementation of large-scale NeuroSoC processors in real-life applications, as is the case for the IBM TrueNorth chip [20] and Intel Loihi [9]. However, they are implemented with digital architectures.
In recent years, several analog CMOS implementations of the LIF neuron have explored the trade‐offs among energy per spike, silicon area, and supply voltage. Indiveri et al. demonstrated a 0.35 µm analog LIF neuron consuming on the order of 900 pJ/spike with a 2 573 $\mu m^{2}$ area at 3.3 V [13], while Wijekoon and Dudek reported 9 pJ/spike in 2 800 $\mu m^{2}$ at 3.3 V in the same technology node [32]. In 65 nm, [37] implemented an Izhikevich neuron based on a ring oscillator achieving 58.5 fJ/spike in 472 $\mu m^{2}$ at 200 mV, and [29] presented a bio-inspired Moris-Lecar design consuming 4 fJ/spike in 35 $\mu m^{2}$ at 200 mV. Ferreira further optimized a streamlined LIF in 55 nm to 3.6 fJ/spike and 85 $\mu m^{2}$ at ±100 mV [10]. These implementations illustrate that scaling and deep sub-threshold operation can dramatically reduce dynamic energy, yet a clear tension remains between achieving both ultra-low energy and minimal area at low voltages.
The primary focus of this work is the design and validation of a scalable Leaky Integrate-and-Fire (LIF) neuron architecture for CMOS implementation. While considerations such as communication protocols (e.g., I2C or SPI) are essential for practical deployments, they fall beyond the scope of this study and will be addressed in future work. Despite rapid advances in digital spiking‐neuroprocessor architectures, there remains an acute need for ultra‐low-power, high-density analog neuron building blocks that can operate at sub-threshold voltages (¡300 mV) and sub-femtojoule energy per event to enable energy-constrained edge AI and biomedical implants. Existing analog LIF implementations either consume more energy or occupy larger areas when scaled to advanced nodes, limiting their applicability in large-scale SNNs on portable and implantable platforms. Our work addresses this gap by presenting the first 28 nm CMOS analog LIF neuron that simultaneously achieves 1.61 fJ/spike, a 34 $\mu m^{2}$ footprint, and 250 mV supply operation. We further validate the design through post‐silicon measurements on 20 fabricated ASICs and develop a hardware-aware software model complete with co-simulation measurement of f-I curves and surrogate-gradient SNN training to demonstrate end-to-end feasibility for edge applications.
@@ II-A LIF Model
This equation represents a balance between two forces: the leaky term $(V_{\text{mem}}(t)-V_{\text{reset}})$, which drives the membrane potential towards the reset potential, and the input current term $R_{m}I_{\text{syn}}(t)$, which pushes the membrane potential in the opposite direction.
@@
  • dtdt: the discrete time interval to update the membrane potential of the model. Smaller values result in more accurate simulations at the cost of increased computational complexity.
@@ II-B LIF Model Computation Sequence
for each time step dtdt do
@@ II-C LIF Implementation in CMOS
The Spike generation block comprises two back-to-back CMOS inverters that generate artificial action potentials at the neuron’s output. The inverters consist of transistors $M_{5}$, $M_{6}$, $M_{7}$, and $M_{8}$. When $V_{mem}$ reaches the inversion threshold $V_{{th}}$, the inverters are activated producing a spike at the output. After producing a spike, the Reset block is activated to discharge the membrane capacitor $C_{mem}$ and prepare the neuron for the next cycle. If the inversion threshold is not reached and no incoming is integrated the leak pathway will slowly discharge the membrane capacitance.
The threshold voltage $V_{th}$ is a critical design parameter that determines when the membrane voltage $V_{mem}$ triggers the spike-generation inverter. A lower $V_{th}$ reduces the required integration time thus increasing the maximum firing frequency. Conversely, a higher $V_{th}$ improves robustness against noise and PVT variations but lengthens the inter-spike interval and marginally increases the average energy per spike due to the extended integration period. In our design, $V_{th}$ is set by the sub-threshold inverter pair M5-M6 and can be expressed as reported in [27]:
$V_{th}=\frac{V_{DD}}{2}+\frac{1}{2}\left(V_{tn}-|V_{tp}|\right)+n_{n}V_{T}\ln\left[\sqrt{\frac{I_{0p}}{I_{0n}}}\right]$ (5)
@@
Leak Pathway
The leak pathway in the proposed circuit is realized implicitly through the subthreshold operation of transistors M3 and M4. These transistors continuously discharge the integration capacitor, introducing a controlled leakage current that emulates the decay dynamics characteristic of the LIF model. This approach allows the circuit to maintain compliance with the LIF model’s fundamental principles without the need for a dedicated leak pathway. The proposed circuit preserves the capability for spatio-temporal coincidence detection by leveraging the interplay between capacitive integration and the leakage current introduced by M3 and M4. Incoming synaptic currents are integrated and compared against the firing threshold within a temporal window determined by the leak dynamics. This enables the neuron to perform temporal filtering, ensuring that coincident inputs are effectively detected.
@@ II-C1 Sub-Threshold Operation
Sub-threshold operation in MOS transistors occurs when the gate-to-source voltage ($V_{GS}$) is below the threshold voltage ($V_{TH}$), resulting in an exponential but weak increase in the drain current ($I_{D}$) with the slight change in $V_{GS}$ [31]. Hence, the drain current requires careful analysis to understand its impact on circuit behavior. In this particular mode, the drain current of a MOS transistor is given by [1, 23]:
$I_{D}\approx I_{D}=I_{0}\frac{W}{L}e^{\frac{V_{GS}-V_{TH}}{nv_{t}}}\left(1-e^{-\frac{V_{DS}}{v_{t}}}\right)$ (6)
Where $n$ and $I_{0}$ are technology dependent and $v_{t}=\frac{k_{B}T}{q}$ is the thermal voltage. What is peculiar in the weak inversion mode is the fact that the drain current is eponentially dependent on the gate-source volatge $V_{GS}$ and the drain-source voltage $V_{DS}$. To better highlight the dependence of the current on $V_{GS}$ and $V_{DS}$ we can rewrite the previous equation based on the analysis from [1]:
$I_{D}=A\cdot e^{\frac{V_{GS}}{nv_{t}}}\cdot\left[e^{\frac{\lambda_{DS}V_{dS}}{nv_{t}}}\left(1-e^{-\frac{V_{DS}}{v_{t}}}\right)\right]$ (7)
Where A represent the stength of the channel inversion on the zero-bias threshold voltage and the junction current going through source/drain to bulk:
$A=I_{0}\frac{W}{L}e^{-\frac{V_{TH0}-\lambda_{BS}V_{BS}}{nv_{t}}}$ (8)
When $V_{DS}$ is relatively high or relatively low compared to $v_{t}$, equation Eq. (9) can be approximated as:
$I_{D}\approx\begin{cases}A\cdot e^{\frac{V_{GS}}{nv_{t}}}\quad\text{ if }V_{DS}\gg v_{t}\\ A\cdot e^{\frac{V_{GS}}{nv_{t}}}\cdot\frac{V_{DS}}{v_{t}}\quad\text{ if }V_{DS}\ll v_{t}\end{cases}$ (9)
From equation Eq. (9), we may rewrite $V_{SG1}$ under the first assumption as:
$V_{SG1}=nv_{t}\cdot\ln{(\frac{I_{syn}}{A})}$ (10)
On the other hand, the drain current $I_{mem}$ in Fig. 1 can be expressed as:
$I_{mem}=A\cdot e^{\frac{V_{GS2}}{nv_{t}}}\;\;\;\text{ if }V_{DS2}\gg v_{t}$ (11)
Given that $V_{SG1}=V_{SG2}$ and under the assumption that $V_{DS}\gg v_{t}$, this yields:
$I_{mem}=A\cdot e^{\frac{nv_{t}\cdot\ln{(\frac{I_{syn}}{A})}}{nv_{t}}}=I_{syn}$ (12)
which simply demonstrate that $I_{syn}$ is copied by $I_{mem}$ and integrated. Another way to interpret the previous simplifications is simply to consider the transistor as a lumped-circuit equivalent. In other words, when looking at the drain and source terminals, the MOS transistor is equivalent to a resistance $R_{syn}$ if $V_{DS}\ll v_{t}$ or a source current $I_{syn}$ if $V_{DS}\gg v_{t}$. Where $R_{syn}$ is the equivalent impedance seen from the drain of $M_{1}$ and given by:
$R_{syn}=\frac{V_{DS}}{I_{D}}=\frac{v_{t}}{A\cdot e^{\frac{V_{GS}}{nv_{t}}}}$ (13)
Given we have $V_{SG1}=V_{DD}-R_{syn}I_{syn}$, we may rewrite $I_{mem}$ as:
$I_{mem}=A\cdot e^{\frac{V_{DD}-R_{syn}I_{syn}}{nv_{t}}}\cdot\frac{V_{DS2}}{v_{t}}\;\;\;\text{ if }V_{DS2}\ll v_{t}$ (14)
Replacin in Eq. 4 yields $V^{max}_{mem}$ peak’s voltage dependence on the input stimulation current $I_{syn}$ as follows:
$V^{max}_{mem}=\frac{A\cdot V_{DS2}}{v_{t}\cdot C_{mem}}e^{\frac{V_{DD}-R_{syn}I_{syn}}{nv_{t}}}+V_{reset}$ (15)
@@
Due to the sub-threshold operation of the transistors, the nominal high output voltage $V_{OH}=V_{DD}-\Delta V_{OH}$ of the CMOS inverter in the spike generation block, comprising cascaded CMOS inverters $M_{5}$-$M_{6}$ and $M_{7}$-$M_{8}$, does not reach $V_{DD}$. The same goes for the nominal low output voltage $V_{OL}=\Delta V_{OL}$. Where both expressions are given by [1, 23]:
$\Delta V_{OH}=v_{t}\frac{A_{n}}{A_{p}}e^{-\frac{V_{DD}}{n_{p}v_{t}}}$ (16)
$\Delta V_{OL}=v_{t}\frac{A_{p}}{A_{n}}e^{-\frac{V_{DD}}{n_{n}v_{t}}}$ (17)
Both expressions yields the resulting voltage swing $V_{OH}-V_{OL}$ that deviates from $V_{DD}$:
$\Delta V$ $=\Delta V_{OH}+\Delta V_{OL}$ (18)
$=v_{t}\left(\frac{A_{n}}{A_{p}}e^{-\frac{V_{DD}}{n_{p}v_{t}}}+\frac{A_{p}}{A_{n}}e^{-\frac{V_{DD}}{n_{n}v_{t}}}\right)$
From $\Delta V_{OH}$ we may conclude that the deviation of $V^{max}_{out}$ from $V_{DD}$ increases exponentially when the supply voltage decreases.